OpenStudy (anonymous):

Use the the limit definition to compute the derivative of the function f(x)=5^x−2 at x=1, and find an equation of the tangent line. f′(1)= At x=1, the tangent line is y=

OpenStudy (aum):

\[ f'(a) = \lim_{h \rightarrow 0}\frac{f(a+h)-f(a)}{h} \\ f'(1) = \lim_{h \rightarrow 0}\frac{f(1+h)-f(1)}{h} = \lim_{h \rightarrow 0}\frac{(5^{1+h}-2)-(5^1-2)}{h} = ?\\ \]

OpenStudy (aum):

\[ f'(1) = \lim_{h \rightarrow 0}\frac{(5^{1+h}-2)-(5^1-2)}{h} = \lim_{h \rightarrow 0}\frac{((5*5^h-2)-(3))}{h} = \\ \lim_{h \rightarrow 0}\frac{5*5^h-5}{h} = \lim_{h \rightarrow 0}\frac{5(5^h-1)}{h} = 5\lim_{h \rightarrow 0}\frac{(5^h-1)}{h} \\ \text{Let } 1/t = 5^h-1 \\ h \rightarrow 0, ~~ t \rightarrow \infty\\ 1+1/t = 5^h \\ \ln(1+1/t) = h\ln(5) \\ h = \ln(1+1/t) / \ln(5) \\ 5\lim_{h \rightarrow 0}\frac{(5^h-1)}{h} = 5\lim_{t \rightarrow \infty}\frac{(1/t)}{\ln(1+1/t)/\ln(5)} = 5\ln(5)\lim_{t \rightarrow \infty}\frac{1}{t\ln(1+1/t)} = \\ 5\ln(5)\lim_{t \rightarrow \infty}\frac{1}{\ln(1+1/t)^t} = 5\ln(5)\frac{1}{\ln(e)} = 5\ln(5)\\ \\ \\ \]

OpenStudy (aum):

when \(x = 1,~~ f(x) = 5^1-2=3\) (1,3) is a point on the curve. \(y - y_1 = m(x - x_1) \\ y - 3 = 5\ln(5)(x - 1) \\ y = 5\ln(5)x + (3 - 5\ln(5)) \) is the equation of the tangent at x = 1.